OpenStudy (anonymous):

Find the excluded value of the rational expression 2x + 6 4x – 8 A. -3 B. -2 C. 0 D. 2

OpenStudy (anonymous):

Did not get.. Excluded value??

OpenStudy (anonymous):

that 's how the problem don't know what else u asking for

OpenStudy (anonymous):

I also don't know what you are asking for... Ha ha ha...

OpenStudy (anonymous):

okay i guess

OpenStudy (anonymous):

Hey, I got it..

OpenStudy (anonymous):

Is this your question : \[\frac{2x + 6}{4x – 8}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I just want to know that is their any divide sign or / sign between 2x + 6 and 4x – 8??

OpenStudy (anonymous):

*there.

OpenStudy (anonymous):

So, excluded value is the value which can make the denominator 0.. As, you know denominator must not be 0, that is why this value is to be excluded and is called Excluded Value..

OpenStudy (anonymous):

So, you have to just put your Denominator = 0: And then find x from it: \[4x - 8 = 0\] Find x here..

OpenStudy (anonymous):

set the denominator = 0, so 4x - 8 = 0 so 4x = 8. Now find the value of x. and this x value will be the excluded value as you cant use the x you find for this equation.

OpenStudy (anonymous):

Can you find the value of x here ???

OpenStudy (anonymous):

@dietrich_harmon we are just expecting a reply from your side..

OpenStudy (anonymous):

yup!!!

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

you can do this @dietrich_harmon.

OpenStudy (anonymous):

Are you able to do this??

OpenStudy (anonymous):

Try once and if any difficulty you will face, then let us know that..